# compat=1.0.0

### v2

@@
Improved constraints on Higgs boson self-couplings with quartic and cubic power dependencies in the cross section
@@
The precise determination of the Higgs boson self-couplings is essential for understanding the mechanism behind electroweak symmetry breaking. However, due to the limited number of Higgs boson pair events at the LHC, only loose constraints have been established so far. Current constraints are based on the assumption that the cross section is a quadratic function of the trilinear Higgs self-coupling within the $\kappa$ framework. Incorporating higher-order quantum corrections from virtual Higgs bosons would significantly alter this function form, introducing new quartic and cubic power dependencies on the trilinear Higgs self-coupling. To derive this new function form, we propose a specialized renormalization procedure that tracks all Higgs self-couplings at each calculation step. Additionally, we introduce renormalization constants for the coupling modifiers within the $\kappa$ framework to ensure the cancellation of all ultraviolet divergences. With the new function forms of the cross sections in both the gluon-gluon fusion and vector boson fusion channels, the upper limit of $\kappa_{\lambda_{\rm 3H}}=\lambda_{\rm 3H}/\lambda_{\rm 3H}^{\rm SM}$ by the ATLAS (CMS) collaboration is reduced from 6.6 (6.49) to 5.4 (5.37). However, extracting a meaningful constraint on the quartic Higgs self-coupling $\lambda_{\rm 4H}$ from Higgs boson pair production data remains challenging. We also present the invariant mass distributions of the Higgs boson pair at different values of $\kappa_{\lambda}$, which could aid in setting optimal cuts for experimental analysis.
@@ 1 Introduction
The current constraints on the trilinear Higgs self-coupling extracted from the Run 2 dataset of Higgs boson pair production at the LHC by the ATLAS and CMS experiments are $-0.6<\kappa_{\lambda_{\rm 3H}}<6.6$ [43] and $-1.24<\kappa_{\lambda_{\rm 3H}}<6.49$ [11], respectively, within the $\kappa$ framework [44], where $\kappa_{\lambda_{\rm 3H}}=\lambda_{\rm 3H}/\lambda_{\rm 3H}^{\rm SM}$ with $\lambda_{\rm 3H}^{\rm SM}$ being the SM value of the trilinear Higgs self-coupling. On the other hand, the process of single Higgs boson production and decay depends on the Higgs self-coupling only via higher-order electro-weak (EW) corrections and can also impose certain constraint [45, 46, 47, 48, 49, 50]. A measurement using the differential fiducial cross section in bins of the Higgs boson transverse momentum sets a constraint of $-5.4<\kappa_{\lambda_{\rm 3H}}<14.9$ [51]. Combined analyses of single- and double-Higgs production result in the constraint, $-0.4<\kappa_{\lambda_{\rm 3H}}<6.3$, at the $95\%$ confidence level, assuming that the new physics changes only the Higgs self-coupling [43].
@@ 2 Renormalization in the $\kappa$ framework
$+\frac{3\lambda_{\rm 3H}^{2}}{16\pi^{2}}\frac{v^{2}}{m_{H}^{2}}(21-4\sqrt{3}\pi)-\frac{9\lambda_{\rm 3H}^{2}v^{2}}{4\pi^{2}}C_{0}[m_{H}^{2},m_{H}^{2},s,m_{H}^{2},m_{H}^{2},m_{H}^{2}]$
@@ 3 Numerical results and improved constraints on the Higgs boson self-coupling
$\delta\sigma^{\kappa_{\lambda}}_{\rm ggF,EW}$ $=(0.075\kappa_{\lambda_{\rm 3H}}^{4}-0.158\kappa_{\lambda_{\rm 3H}}^{3}-0.006\kappa_{\lambda_{\rm 3H}}^{2}\kappa_{\lambda_{\rm 4H}}-0.058\kappa_{\lambda_{\rm 3H}}^{2}$
@@
$\delta\sigma_{\rm VBF,EW}^{\kappa_{\lambda}}$ $=(0.0215\kappa_{\lambda_{\rm 3H}}^{4}-0.0324\kappa_{\lambda_{\rm 3H}}^{3}-0.0019\kappa_{\lambda_{\rm 3H}}^{2}\kappa_{\lambda_{\rm 4H}}-0.0043\kappa_{\lambda_{\rm 3H}}^{2}$
$~{}~{}~{}+0.0151\kappa_{\lambda_{\rm 3H}}\kappa_{\lambda_{\rm 4H}}-0.0211\kappa_{\lambda_{\rm 4H}})~{}~{}{\rm fb}$ (13)
for the VBF channel. We have computed all the $\mathcal{O}(\lambda_{\rm 3H}^{i}),i\geq 2$ contributions in the amplitude. The above $\kappa_{\lambda_{\rm 3H}}^{2}$ terms arise because we want to keep the cancellation relation between the $\mathcal{O}(\lambda_{\rm 3H})$ and $\mathcal{O}(1)$ amplitudes at LO. The cubic $\kappa_{\lambda_{\rm 3H}}^{3}$ and quartic $\kappa_{\lambda_{\rm 3H}}^{4}$ terms appear for the first time up to this perturbative order. Though their coefficients are rather small, they provide notable corrections to the cross section if $\kappa_{\lambda_{\rm 3H}}$ is chosen much larger than 1. As seen from Table 1, the $\lambda$ dependent corrections in the ggF (VBF) channel reach $91\%$ ($82\%$) of the LO cross section for $\kappa_{\lambda_{\rm 3H}}=6$.
@@
1 1 16.7 31.2 -0.225 1.71 1.69 $-2.30\times 10^{-2}$
3 1 8.59 18.4 1.28 3.59 3.53 $8.35\times 10^{-1}$
6 1 67.3 161 60.6 25.1 24.6 20.7
1 3 16.7 31.2 -0.393 1.71 1.69 $-3.89\times 10^{-2}$
1 6 16.7 31.2 -0.646 1.71 1.69 $-6.27\times 10^{-2}$
3 3 8.59 18.4 1.30 3.59 3.53 $8.50\times 10^{-1}$
6 6 67.3 161 61.0 25.1 24.6 20.7
@@
In addition, there is a new dependence on the quartic Higgs self-coupling $\lambda_{\rm 4H}$. Because this dependence is only linear and the corresponding coefficients are small, their contributions are negligible. From Table 1, it can be observed that the cross section varies by $0.6\%$ when $\kappa_{\lambda_{\rm 4H}}$ changes from 1 to 6 while keeping $\kappa_{\lambda_{\rm 3H}}=6$. As a consequence, we do not expect that a meaningful constraint on the quartic Higgs self-coupling can be extracted from Higgs boson pair production at the LHC.
@@
In Fig. 3, we show different perturbative predictions for the cross sections of both ggF and VBF Higgs boson pair production at the 13 TeV LHC as a function of $\kappa_{\lambda_{\rm 3H}}=\kappa_{\lambda_{\rm 4H}}=\kappa_{\lambda}$. It is evident that higher-order perturbative corrections dramatically change the functional form. The current experimental upper limit by the ATLAS (CMS) collaboration on $\kappa_{\lambda}$ is $6.6$ (6.49) based on the theoretical predictions at QCD NNLO in the ggF channel and NNNLO in the VBF channel; see Table 1. Taking $\delta\sigma^{\kappa_{\lambda}}_{\rm EW}$ corrections into account and assuming that the QCD and EW corrections are factorizable, the upper limit is narrowed down to 5.4 (5.37). These limits are almost the same when keeping $\kappa_{\lambda_{\rm 4H}}=1$. If the scale uncertainties are considered [74], the upper limit spans in the range $(6.5,6.8)$ in the ATLAS result, which would decrease to $(5.4,5.6)$ after including higher power dependence. In the CMS result, the upper limit changes from (6.40, 6.67) to (5.31, 5.48). The lower limits are only modified slightly.
@@ 4 Conclusion
We propose a proper renormalization procedure to explicitly retain the Higgs self-couplings at each calculation step and introduce renormalization of the coupling modifiers to ensure the cancellation of all ultraviolet divergences. We present numerical results of the cross sections of both the ggF and VBF channels at the LHC including higher power dependencies on the Higgs self-coupling. With these improved function forms, we demonstrate that the upper limit set by the ATLAS (CMS) collaboration on the trilinear Higgs self-coupling normalized to its SM value is reduced from 6.6 (6.49) to 5.4 (5.37). This more precise constraint is achieved without more data being analyzed, underscoring the critical importance of incorporating higher power dependencies on the Higgs self-coupling in the cross section.

### v3

@@
Improved constraints on Higgs boson self-couplings with quartic and cubic power dependencies of the cross section
@@
Precise determination of the Higgs boson self-couplings is essential for understanding the mechanism underlying electroweak symmetry breaking. However, owing to the limited number of Higgs boson pair events at the LHC, only loose constraints have been established to date. Current constraints are based on the assumption that the cross section is a quadratic function of the trilinear Higgs self-coupling within the $\kappa$ framework. Incorporating higher-order quantum corrections from virtual Higgs bosons would significantly alter this functional form, introducing new quartic and cubic power dependencies on the trilinear Higgs self-coupling. To derive this new functional form, we propose a specialized renormalization procedure that tracks all Higgs self-couplings at each calculation step. Additionally, we introduce renormalization constants for coupling modifiers within the $\kappa$ framework to ensure the cancellation of all ultraviolet divergences. With new functional forms of the cross sections in both the gluon-gluon fusion and vector boson fusion channels, the upper limit of $\kappa_{\lambda_{\rm 3H}}=\lambda_{\rm 3H}/\lambda_{\rm 3H}^{\rm SM}$ set by the ATLAS (CMS) collaboration is reduced from 6.6 (6.49) to 5.5 (5.39). However, extracting a meaningful constraint on the quartic Higgs self-coupling $\lambda_{\rm 4H}$ from Higgs boson pair production data remains challenging. We also present the invariant mass distributions of the Higgs boson pair at different values of $\kappa_{\lambda}$, which could assist in setting optimal cuts for experimental analysis.
@@ 1 Introduction
[…]
@@ 2 Renormalization in the $\kappa$ framework
$+\frac{3\lambda_{\rm 3H}^{2}}{8\pi^{2}}\frac{v^{2}}{m_{H}^{2}}\left(12-\frac{7\pi}{\sqrt{3}}\right)-\frac{9\lambda_{\rm 3H}^{2}v^{2}}{4\pi^{2}}C_{0}[m_{H}^{2},m_{H}^{2},s,m_{H}^{2},m_{H}^{2},m_{H}^{2}]$
@@ 3 Numerical results and improved constraints on the Higgs boson self-coupling
$\delta\sigma^{\kappa_{\lambda}}_{\rm ggF,EW}$ $=(0.067\kappa_{\lambda_{\rm 3H}}^{4}-0.105\kappa_{\lambda_{\rm 3H}}^{3}-0.006\kappa_{\lambda_{\rm 3H}}^{2}\kappa_{\lambda_{\rm 4H}}-0.166\kappa_{\lambda_{\rm 3H}}^{2}$
@@
$\delta\sigma_{\rm VBF,EW}^{\kappa_{\lambda}}$ $=(0.0196\kappa_{\lambda_{\rm 3H}}^{4}-0.0235\kappa_{\lambda_{\rm 3H}}^{3}-0.0014\kappa_{\lambda_{\rm 3H}}^{2}\kappa_{\lambda_{\rm 4H}}-0.0165\kappa_{\lambda_{\rm 3H}}^{2}$
$~~~+0.0126\kappa_{\lambda_{\rm 3H}}\kappa_{\lambda_{\rm 4H}}-0.0193\kappa_{\lambda_{\rm 4H}})~~{\rm fb}$ (13)
[…]
@@
1 1 16.7 31.2 -0.287 1.71 1.69 $-2.85\times 10^{-2}$
3 1 8.59 18.4 1.15 3.60 3.53 $8.09\times 10^{-1}$
6 1 67.3 161 58.7 25.1 24.6 19.7
1 3 16.7 31.2 -0.455 1.71 1.69 $-4.47\times 10^{-2}$
1 6 16.7 31.2 -0.708 1.71 1.69 $-6.91\times 10^{-2}$
3 3 8.59 18.4 1.17 3.60 3.53 $8.21\times 10^{-1}$
6 6 67.3 161 59.1 25.1 24.6 19.7
@@
[…]
@@
In Fig. 3, we show different perturbative predictions for the cross sections of both ggF and VBF Higgs boson pair production at the 13 TeV LHC as a function of $\kappa_{\lambda_{\rm 3H}}=\kappa_{\lambda_{\rm 4H}}=\kappa_{\lambda}$. It is evident that higher-order perturbative corrections dramatically change the functional form. The current experimental upper limit by the ATLAS (CMS) collaboration on $\kappa_{\lambda}$ is 6.66.66.6 (6.49) based on the theoretical predictions at QCD NNLO in the ggF channel and NNNLO in the VBF channel; see Table 1. Taking $\delta\sigma^{\kappa_{\lambda}}_{\rm EW}$ corrections into account and assuming that the QCD and EW corrections are factorizable, the upper limit is narrowed down to 5.5 (5.39). These limits are almost the same when keeping $\kappa_{\lambda_{\rm 4H}}=1$. If the scale uncertainties are considered [74], the upper limit spans in the range (6.5,6.8)(6.5,6.8)( 6.5 , 6.8 ) in the ATLAS result, which would decrease to (5.4,5.6)(5.4,5.6)( 5.4 , 5.6 ) after including higher power dependence. In the CMS result, the upper limit changes from (6.40, 6.67) to (5.33, 5.50). The lower limits are only modified slightly.
@@ 4 Conclusion
We propose a proper renormalization procedure to explicitly retain the Higgs self-couplings at each calculation step and introduce renormalization of the coupling modifiers to ensure the cancellation of all ultraviolet divergences. We present numerical results of the cross sections of both the ggF and VBF channels at the LHC including higher power dependencies on the Higgs self-coupling. With these improved function forms, we demonstrate that the upper limit set by the ATLAS (CMS) collaboration on the trilinear Higgs self-coupling normalized to its SM value is reduced from 6.6 (6.49) to 5.5 (5.39). This more precise constraint is achieved without more data being analyzed, underscoring the critical importance of incorporating higher power dependencies on the Higgs self-coupling in the cross section.